# Kinetic Mixing, Proton Decay and Gravitational Waves in SO(10)

### v1

@@ 1 Introduction
The quest for realistic grand unification based on $SO(10)$ [1, 2] predicts topologically stable magnetic monopoles [3] and proton decay [4, 5], and they are yet to be observed [6, 7]. In addition, topologically stable strings [8], and composite structures such as walls bounded by strings (WBS) [9, 10, 11] and monopoles connected by strings [12] appear in $SO(10)$ depending on the symmetry breaking patterns. The latter composite structures can give rise to the formation of the quasistable string (QSS) network [13, 14]. The recent NANOGrav 15 year data [15, 16, 17] and other pulsar timing array (PTA) data [18, 19, 20, 21] impose a lower bound of $G\mu\lesssim 10^{-10}$ on topologically stable cosmic strings, where $\mu$ denotes the mass per unit length of the string. However, the PTA data is compatible with the gravitational wave background emitted by QSS and WBS [11, 22] networks with superheavy strings having $G\mu\sim 10^{-6}$. Another example is based on the so-called metastable strings (MSS) [23, 24], where the topologically unstable strings decay via quantum mechanical tunneling. For some recent studies of MSS based on grand unification see Refs. [25, 26, 27, 28, 29, 30, 31].
@@ 5 Gravitational waves from quasistable strings
Figure 3: Gravitational wave background from quasistable cosmic strings with $G\mu=10^{-6}$ and $t_{M}=10^{2.5}$ sec. The red violins depict the posteriors of an HD correlated free spectral reconstruction of the NANOGrav 15 year data. We assume the string network enters a scaling regime at the time $t_{F}=10^{-10}$ sec after their horizon re-entry in order to satisfy the LIGO-VIRGO third run data. The gray region depicts the bound from Big Bang Nucleosynthesis (BBN) [75]. We have shown the power-law integrated sensitivity curves [76, 77] for planned experiments, namely, HLVK [78], CE [79], ET [80], DECIGO [81], BBO [82, 83], LISA [84, 85], $\mu$Ares [86] and SKA [87, 88].
@@
where $\frac{d^{2}R}{dz\,dl}$ is the burst rate coming from a loop distribution $n(l,t)$, $d_{H}$ denotes the particle horizon at time $t(z)$, and $H_{0}$ is the Hubble parameter at the present era. Extensive studies on the gravitational waves from strings can be found in the literature [92, 93, 94, 95, 96, 97, 98, 99, 90, 100, 101, 102, 103, 24].
@@
Figure 3 depicts the gravitational wave background emitted by the quasistable cosmic strings with $G\mu=10^{-6}$ and a monopole re-entry time $t_{M}=10^{2.5}$ sec. In order to satisfy the advanced LIGO-VIRGO bound, we assume that the string network enters a scaling regime at a late time $t_{F}=10^{-10}$ sec after their horizon re-entry. We utilize the wrapper PTArcade [104] to reconstruct the Bayesian credible region of the model parameters $G\mu$ and $t_{M}$ in QSS with the NANOGrav 15 year data. We employ the Ceffyl package [105] with the Hellings-Downs (HD) correlation [106] between the pulsars to obtain the posterior distribution. The triangular plot in Figure 4 shows the posterior distribution of $G\mu$ and $t_{M}$. Table 1 shows the 68% and 95% Bayesian credible intervals of the model parameters.

### v2

@@ 1 Introduction
The quest for realistic grand unification based on $SO(10)$ [1, 2] predicts topologically stable magnetic monopoles [3] and proton decay [4, 5], and they are yet to be observed [6, 7]. In addition, topologically stable strings [8], and composite structures such as walls bounded by strings (WBS) [9, 10, 11] and monopoles connected by strings [12] appear in $SO(10)$ depending on the symmetry breaking patterns. The latter composite structures can give rise to the formation of the quasistable string (QSS) network [13, 14]. The recent NANOGrav 15 year data [15, 16, 17] and other pulsar timing array (PTA) data [18, 19, 20, 21] impose a lower bound of $G\mu\lesssim 10^{-10}$ on topologically stable cosmic strings, where $\mu$ denotes the mass per unit length of the string. However, the PTA data is compatible with the gravitational wave background emitted by QSS and WBS [11, 22] networks with superheavy strings having $G\mu\sim 10^{-6}$. Another example is based on the so-called metastable strings (MSS) [23, 24], where the topologically unstable strings decay via quantum mechanical tunneling. For some recent studies of MSS based on grand unification see Refs. [25, 26, 27, 28, 29, 30, 31]. Stable string networks with a reduced intercommutation probability ($\mathcal{P}$) such as cosmic superstrings are expected to generate stochastic gravitational waves enhanced by a factor $\sim 1/\mathcal{P}$ [32, 33, 34, 35, 36, 37, 38, 39]. In turn, they can explain the PTA data with $G\mu\sim 10^{-12}-10^{-11}$ and $\mathcal{P}\sim 10^{-3}-10^{-2}$ [17, 40, 41].
@@ 5 Gravitational waves from quasistable strings
Figure 3: Gravitational wave background from quasistable cosmic strings with $G\mu=10^{-6}$ and $t_{M}=10^{2.5}$ sec. The red violins depict the posteriors of an HD correlated free spectral reconstruction of the NANOGrav 15 year data. We assume the string network enters a scaling regime at the time $t_{F}=10^{-10}$ sec after their horizon re-entry in order to satisfy the LIGO-VIRGO third run data [85]. The gray region depicts the bound from Big Bang Nucleosynthesis (BBN) [86]. We have shown the power-law integrated sensitivity curves [87, 88] for planned experiments, namely, HLVK [89], CE [90], ET [91], DECIGO [92], BBO [93, 94], LISA [95, 96], $\mu$Ares [97] and SKA [98, 99].
@@
where $\frac{d^{2}R}{dz\,dl}$ is the burst rate coming from a loop distribution $n(l,t)$, $d_{H}$ denotes the particle horizon at time $t(z)$, and $H_{0}$ is the Hubble parameter at the present era. Extensive studies on the gravitational waves from strings can be found in the literature [103, 104, 105, 106, 107, 108, 109, 110, 101, 111, 112, 113, 114, 24, 115].
@@
Figure 3 depicts the gravitational wave background emitted by the quasistable cosmic strings with $G\mu=10^{-6}$ and a monopole re-entry time $t_{M}=10^{2.5}$ sec. In order to satisfy the advanced LIGO-VIRGO bound [85], we assume that the string network enters a scaling regime at a late time $t_{F}=10^{-10}$ sec after their horizon re-entry. We utilize the wrapper PTArcade [116] to reconstruct the Bayesian credible region of the model parameters $G\mu$ and $t_{M}$ in QSS with the NANOGrav 15 year data. We employ the Ceffyl package [117] with the Hellings-Downs (HD) correlation [118] between the pulsars to obtain the posterior distribution. The triangular plot in Figure 4 shows the posterior distribution of $G\mu$ and $t_{M}$. Table 1 shows the 68% and 95% Bayesian credible intervals of the model parameters.